OpenStudy (anonymous):

Solve: y/2-1/8=3y-5/8

OpenStudy (anonymous):

y=1/5

OpenStudy (anonymous):

y/2-1/8 = 3y-5/8 (4y-1)/8 = (24y-5)/8 (common denominators) 4y-1 = 24y-5 -1+5 = 24y-4y (collect like terms together) 4 = 20y y = 1/5